40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Wed Oct 20 11:31:06 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever 
    Reply  Tue Oct 26 10:01:54 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever TiltedCantilever.pdfcantileversurface.pdfclampsurface.pdf
       Reply  Thu Nov 18 11:09:24 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever CavitySetup.pdfGeometry01.jpeg
          Reply  Wed Dec 1 13:21:54 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever IMG_1250.pngIMG_1252.pngIMG_1251.pngIMG_1253.pngIMG_1254.png
             Reply  Thu Dec 2 13:40:57 2021, shruti, Update, PSOMA, Progress and cavity parameters/ lock parameters 
Message ID: 2829     Entry time: Wed Dec 1 13:21:54 2021     In reply to: 2828     Reply to this: 2830
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: Towards locking PSOMA with cantilever 

In order to improve the acoustic isolation of the box I used rubber tubing cut as depicted in Attachment 1 and attached to the edge of the plastic box (as seen in Attachment 2).

Attachment 2 also shows one of the four dog clamps I used to secure the box. Finally I placed an oscilloscope on top of the box to provide some weight.

The box has 4 holes for the input, reflected, transmitted-to-PD, and transmitted-to-camera beams.

 

PDH signal

In elog 2731, the PDH signal is almost 1 V pk-pk, but recently I had only been seeing a very noisy signal ~10 mV pk-pk (purple trace in Attachment 4. Transmitted light is yellow, reflected light is blue). I adjusted the lengths and removed the attenuators in the path of the EOM but it did not seem to change.

While going through the path of the cables once more, I realized that signal from the OCXO was being sent to the wrong port on the rack mount (Attachment 5). It was earlier on D3, and I later changed it to A4 which was labelled 'S EOM', but the actual location was D4 which I identified from this diagram. Now I see a reasonably large PDH signal as expected with ~800mV pk-pk (video here).

Attachment 1: IMG_1250.png  12.204 MB  | Hide | Hide all
IMG_1250.png
Attachment 2: IMG_1252.png  11.022 MB  | Hide | Hide all
IMG_1252.png
Attachment 3: IMG_1251.png  15.975 MB  | Hide | Hide all
IMG_1251.png
Attachment 4: IMG_1253.png  16.465 MB  Uploaded Wed Dec 1 17:07:32 2021  | Hide | Hide all
IMG_1253.png
Attachment 5: IMG_1254.png  10.311 MB  Uploaded Wed Dec 1 17:08:12 2021  | Hide | Hide all
IMG_1254.png
ELOG V3.1.3-